Assume again that last year Brock had an EBIT of $1600, Depreciation of 600, capital expenditures of 700, operating NWC of 1000 (this is not change in NWC)and a tax rate of .25. However, Allan’s management believes that Brock will grow at 18% for the next 3 years, and at a constant rate of 5% thereafter. Recompute the value of Brock, assuming that its WACC will remain 10%.
Answer #1

FCFF = EBIT(1 - t) + Depreciation - Change in Capex - Change in Net Working Capital

FCFF = 1,600(1 - 0.25) + 600 - 700 - 0

FCFF0 = $1,100

Value of Firm = 1,100(1.18/1.10) + 1,100(1.18/1.10)2 + 1,100(1.18/1.10)3 + 1,100(1.18)3(1.05)/(0.10 - 0.05)(1.10)3

Value of Firm = 1,180 + 1,265.82 + 1,357.88 + 28,515.43

Value of Firm = $32,319.13